Apple Company owns 5% (50 of the 1000 shares) of the stock in Banana Company which it purchased for $5000. Apple has no influence over Banana Company. On January 1st Banana Company has a two for one split. On February 3rd Apple sold 10 of its shares in Banana Company for $300. What gain or loss does Apple show on its sale of Banana Company?

Select an answer and submit. For keyboard navigation, use the up/down arrow keys to select an answer.

a

Gain $50

b

Loss $200

c

Loss $4700

d

Gain 300

e

No gain or loss on partial sale
Answer #1

Ans. (B). loss of $200.

Cost of 50 shares = $5000

cost of individual share = $100

After split of shares, cost of each share = 100 / 2 = $ 50 each

Hence, Cost of 10 shares = 50 * 10 = 500

Acctual sales price = $ 300

Thus , loss = $ 200.